OpenStudy (anonymous):

for every serving off potatoes Manny boils 1 1/2 potatoes . How many servings of potatoes can be made from 15 potatoes ?

OpenStudy (anonymous):

please help

myininaya (myininaya):

1serving =3/2 potatoes we need 15 potatoes on the left hand side what can we multiply 3/2 by that will give us 15 you could solve this equation to figure out 3/2*x=15 x=15*2/3=5*2=10 so if we multiply both sides of 1serving=3/2 potatoes by 10 we get 10servings=10*3/2potatoes 10servings=15potatoes

OpenStudy (karatechopper):

10 servings